A rectangular parking lot is 110 yards and 45 yards wide. it costs about $.60 to pave each square foot of the parking lot with asphalt. about how much will it cost to pave the parking lot ?
Found 2 solutions by nyc_function, Alan3354:
Answer by nyc_function(2741)

A = L*W
A = 110*45
A = 4950 yards
There are 3 feet per yards.
So, 4950/3 = 1650 feet.
Now, 0.60 times 1650 = $990

Answer by Alan3354(69443)

A = L*W
A = 110*3 ft*45*3 ft
A = 44550 sq ft
$0.60 times 44550 = $26730